OpenStudy (anonymous):

Imagine two solutions with the same concentration and the same boiling point, but one has ethanol as the solvent and the other has carbon tetrachloride as the solvent. Determine that concentration and boiling point.

OpenStudy (shiraz14):

Kb for ethanol = 1.22; normal b.p. for ethanol = 78.4 deg C Kb for CCl4 = 5.03; normal b.p. for CCl4 = 76.8 deg C Since the b.p. for ethanol & CCl4 solutions in this context are the same, 78.4 + 1.22 m = 76.8 + 5.03 m Concentration of solution (m) = (78.4-76.8)/(5.03-1.22) = 0.41995 M b.p. of solution = 78.4 + (1.22*0.41995) = 78.912 deg C

OpenStudy (anonymous):

Thank you very much!

OpenStudy (shiraz14):

You're welcome.